40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4996     Entry time: Wed Jul 20 06:52:01 2011
Author: Suresh 
Type: Update 
Category: LSC 
Subject: ABSL - PSL beat lock 

The ABSL locking setup to the PSL is down. 

According the plan, I started to use the IR beam dumped after the doubling crystal for the IR beat lock (Sonali's project).  The beat lock was disturbed when I shifted some clamps to make way for a few mirrors.  So I set about fixing the beat lock.  I reobtained the lock but noticed that the net beam power reaching the Newfocus 1611 detector was around 15mW.  10mW from the ABSL and 5mW from PSL.

This is much too high as the maximum allowed on 1611 is 2mW. 

I therefore started to adjust the power levels by using  Y1-1064-45S mirrors at non-45 deg angles.  However Rana pointed out that this would lead to amplitude noise due to the mirror vibrations.  I then switched to using beam splitters as pick offs.   This is better than using neutral density filters since the back scatter is lower this way.

David wanted some of the ABSL beam for his SURF student.  So I changed the mirror after beam expanding telescope on the ABSL route to provide this power.  We also installed a pair of half wave plates and a PBS to allow us smooth power level control on this beam.

The beat lock setup is now down and needs to be completed for PRCL and SRCL measurements.

 

 

ELOG V3.1.3-